OpenStudy (anonymous):

If a rancher sold 1/3 of his cattle and then sold another 1/4 of the original amount, what part of his cattle was sold? A. 1/7 B. 2/12 C. 7/12

OpenStudy (anonymous):

7/12 I believe.

OpenStudy (anonymous):

because 1/3 + 1/4 = 7/12

OpenStudy (anonymous):

In order to solve this problem, we have to make each one of the fractions have a common denominator. So, take 1/3 and 1/4 and find out the common denominator. The easiest common denominator to deal with is 12. So, 4 goes into 12 3 times, so multiply the top and bottom of the 1/4 fraction. You will now have 3/12. Do the same thing with 1/3. 3 goes into 12 4 times so multiply the entire fraction 1/3 by 4. You will now have 4/12. We want to know how much of his cattle was sold, so we add. 4/12 + 3/12=7/12 (Just add the numerator and put it over the common denominator!).